40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 766     Entry time: Wed Jul 30 13:08:44 2008
Author: Max Jones 
Type: Update 
Category: Computer Scripts / Programs 
Subject: Weekly Summary 
This week I've been working on the noise budget script. The goal is to add Siesmic, Darm, Mich, Prc and magnetometer noise. I believe I've added Seismic noise in a reasonable and 40m specific manner (please see the attached graph). The seismic noise in the noise budget at 100 Hz was 10 times higher than that predicted by Rana in elog #718. This could be due to the fact that graph is taken from data today when the device is unlocked and construction workers are busy next door. I am currently trying to fix the getDarm.m file to add the DARM source to the noise budget. I have run into several problems, the most pressing of which was that the C1:LSC-DARM_ERR channel is zero except with the interferometer is being locked. According to Rob, we only save data for approximately a day (we save trends for much longer but this is insufficient for the noise budget script) and sometimes we are not locked the night before. Rob showed me how I may introduce an artificial noise in the DARM_ERR signal but I'm having trouble making the script output a graphic. I'm still unsure how to make the getDarm function 40m specific.

Today I will start working on my second progress report and abstract.
Attachment 1: C1_NoiseBudgetPlot.pdf  71 kB  | Hide | Hide all
C1_NoiseBudgetPlot.pdf
ELOG V3.1.3-